# DistTrain: Addressing Model and Data Heterogeneity with Disaggregated Training for Multimodal Large Language Models

### v2

@@ 8.1 Overall Performance
Setup. We first compare the overall performance of DistTrain against Megatron-LM on a large-scale GPU cluster (up to 1296 GPUs). We retrofit Megatron-LM to support multimodal LLM training by integrating modality encoder and generator into the training pipeline. Megatron-LM employs rigid model orchestration as described in §2.1. In Megatron-LM, we set the PP size of the LLM backbone to 1, 2, and 10 for Llama3-7B, Llama3-13B, and Llama3-70B. PP size is set to 1 for modality encoder and generator. TP size is set to 8. As for DistTrain, the parallelism strategy is determined by disaggregated model orchestration. In our experiments, one GPU is able to facilitate training ViT and Stable-Diffusion. We replicate the encoder and generator across the GPUs within the TP group to process different images, whereas TP itself is not used. We set global batch size to 1920.
The experimental results are shown in Figure 15 and Figure 16. Figure 15 shows the MFU. Figure 16 shows the training throughput and marks the number of GPUs used in each experiment. The different GPU numbers are due to the varying model orchestration strategies of DistTrain and Megatron-LM. We summarize the experiment as follows.
@@
As shown in Figure 15, DistTrain achieves 51.8%-54.7% MFU in large-scale multimodal LLM training. This performance closely approximates that of state-of-the-art unimodal (i.e., text) LLM training [29], which demonstrates the effectiveness of DistTrain in addressing the model and data heterogeneity in multimodal LLM training.
@@
DistTrain significantly outperforms Megatron-LM, delivering 1.7-2.8 the MFU and 1.7-2.2 the training throughput when training MLLM-9B and MLLM-15B with a similar number of GPUs. These performance gains largely stem from DistTrain’s disaggregated model orchestration. Megatron-LM’s rigid strategy often leads to GPU underutilization, since it assigns too many GPUs to the modality encoder and generator. In contrast, DistTrain adaptively adjusts model orchestration based on specific model and data demands. Additionally, DistTrain’s disaggregated data reordering technique further boosts efficiency.
@@
In MLLM-72B training scenario, DistTrain also outperforms Megatron-LM by 1.2 on MFU and 1.3 on training throughput with a similar number of GPUs. The high image resolution prolongs the execution time of the multimodal module, which introduces pipeline bubbles in LLM backbone. DistTrain addresses this by allocating additional GPUs to these modules to balance the pipeline. The disaggregated data reordering strategy continues to diminish data heterogeneity, thereby increasing training efficiency.
8.2 Deep Dive into DistTrain
@@
In this subsection, we perform microbenchmarks to evaluate the effectiveness of each DistTrain’s component. Our experiments utilize up to 98 NVIDIA A800 GPUs. We set the global batch size to 128, 64, and 40 for MLLM-9B, MLLM-15B, and MLLM-72B, respectively.
8.2.1 Disaggregated Model Orchestration
@@
We measure the MFU and training throughput of DistTrain and other model orchestration strategies, including a comparison with Megatron-LM and a ratio-based approach. The ratio-based strategy allocates GPUs according to the computational demands (flops) of each module. We also annotate the total number of GPUs utilized in the throughput chart. The experimental results are shown in Figure 17. DistTrain consistently outperforms the baseline strategies, achieving 1.3-2.7 higher MFU and 1.4-2.7 higher training throughput. Although the ratio-based strategy outperforms Megatron-LM’s rigid strategy, it still lags behind DistTrain since it neglects the intricate performance model (§5.1) of multimodal LLM training. DistTrain’s disaggregated model orchestration optimally balances computational loads across the three modules and achieves high resource utilization.
@@
We also evaluate the running time of DistTrain’s disaggregated model orchestration under different training settings, as detailed in Table 2. The algorithm completes in under one second. The overhead is negligible compared to the days or even weeks required for overall training.
8.2.2 Disaggregated Data Reordering
We evaluate the effectiveness of DistTrain’s disaggregated data reordering by comparing it against the random order, while keeping other components the same. The effectiveness is gauged through metrics such as MFU and training throughput (TPT.). We use the optimal resource allocation and parallelism strategy decided by DistTrain’s disaggregated model orchestration. Given that the model orchestration strategy remains unchanged, the number of GPUs is not shown. The experimental settings are the same as those in §8.2.1. The results are shown in Figure 18. DistTrain consistently outperforms the baseline, achieving 1.03-1.11 higher MFU and training throughput. The performance gap becomes more pronounced as the model size decreases. This is because the smaller model size leads to a higher data parallelism (DP) size, which causes more inter-microbatch heterogeneity. In essence, DistTrain’s disaggregated data reordering effectively mitigates data heterogeneity and enhances the training efficiency. We do not measure the running time of the reordering algorithm as it operates on dedicated CPU nodes asynchronously. It does not interfere with the GPU training process. The only overhead is the network delay which is evaluated in §8.3.
@@ 8.2.2 Disaggregated Data Reordering
8.3 Case Study
In this subsection, we first evaluate DistTrain under different frozen training settings. We then evaluate the overhead of CPU preprocessing disaggregation and the effectiveness of StepCCL in mitigating the TP overhead.
8.3.1 Frozen Training
We conduct a frozen training experiment under four specific training settings: complete module freezing, exclusive encoder training, exclusive LLM training, and exclusive generator training. We keep training the projectors. In these scenarios, frozen modules neither compute weight gradients nor update weights. All other experimental setup aligns with those detailed in §8.2. We evaluate the MFU and training throughput (TPT.) of DistTrain compared to Megatron-LM. The experimental results are presented in Figure 19 and Figure 20. DistTrain consistently outperforms Megatron-LM across all frozen training configurations, achieving 1.4-2.9 higher MFU and 1.2-2.9 higher training throughput. This pronounced performance gap underscores the challenges posed by Megatron-LM’s rigid model orchestration in complex training environments. In contrast, DistTrain adaptively adjusts model orchestration based on training settings and consistently achieves high resource utilization.
@@ 8.3.1 Frozen Training
(a) All modules freezing. (b) Exclusive encoder training. (c) Exclusive LLM training. (d) Exclusive generator training.
(a) All modules freezing. (b) Exclusive encoder training. (c) Exclusive LLM training. (d) Exclusive generator training.
@@
Overhead of data preprocessing. We conduct an experiment to evaluate the overhead of data preprocessing, including decompression and reordering. Setting the DP size to one, we measure the average data preprocessing time per iteration on the GPU training side. We then compare data preprocessing time with and without CPU preprocessing disaggregation and use varying numbers of images and different image resolutions for one training iteration. The results, depicted in Figure 21, indicate that the disaggregation significantly reduces preprocessing time from seconds to milliseconds. The first parameter in the x-axis represents the number of images, while the second parameter denotes the image resolution. In production training (§8.1), iteration times range from seconds to tens of seconds. Preprocessing overhead, initially counted in seconds, significantly interferes with training. With disaggregated data preprocessing, the overhead reduces to milliseconds, which is negligible relative to total iteration time.
Mitigating TP overhead with StepCCL. To evaluate the effectiveness of StepCCL in mitigating the TP overhead, we conduct an experiment that measures the iteration time of the LLM backbone with training of one single PP stage (i.e., one minimal TP group) under various TP sizes. We compare the iteration time with and without StepCCL enabled. The results are shown in Figure 22. StepCCL significantly reduces the iteration time by overlapping the TP communication with computation. It outperforms the baseline by 1.1-1.17 when the TP size is 4 and 1.15-1.17 when the TP size is 8. The gains are more pronounced at large TP size, where communication overhead is more substantial. These findings confirm that StepCCL effectively mitigates TP overhead.
9 Discussion
@@
Parallelism optimization. Many studies [47, 48, 49] refine parallelism strategies for deep learning models, but they fall short for multimodal LLMs due to their expansive search spaces and large cluster requirements. These methods separately optimize different strategies, resulting in inefficient parallelism Additionally, these methods generally assume data homogeneity, standardizing computations across all training samples. This assumption does not hold for multimodal LLMs due to the data heterogeneity in modality inputs of each sample. In contrast, DistTrain leverages the specific training pattern of multimodal LLMs, creating a customized model orchestration problem that integrates tensor, pipeline, and data parallelism simultaneously. Besides, DistTrain leverages disaggregated data reordering to harness the data heterogeneity.
@@
Sequence and expert parallelism. Sequence parallelism (SP) [50] is designed to partition the training sequence into multiple subsequences for parallel training. It addresses the challenges of processing long sequences in LLMs. Expert parallelism (EP) [51], specifically devised for mixture-of-experts (MoE) LLMs [52], enables parallel training of multiple feedforward network (FFN) experts. These parallelism strategies are orthogonal to multimodal LLM training. In DistTrain, both SP and EP are integrated into the training framework. DistTrain treats SP and EP sizes as predefined parameters in the disaggregated model orchestration optimization problem.
@@
10 Related Work
LLM training. Many efforts have been made to optimize the training of LLMs from system perspectives. For LLM pretrain, Megatron-LM [17] and DeepSpeed-Megatron [53] propose customized 3D-parallelism and are de facto standards for training large LLMs. DeepSpeed-ZeRO [32] and Pytorch-FSDP [54] reduce redundant memory consumption in data parallelism. HPN [55] proposes a new dual-ToR network architecture trailed for LLM training to reduce ECMP conflict. A set of works [56, 29, 42, 43, 44] overlap the communication and computation operators in LLM training. Fault tolerance through replication and checkpoint is advanced in large training clusters by studies [29, 57]. Efforts like [58, 59, 60] further optimize recovery process in cloud spot instance scenarios. These system optimizations of LLM training are orthogonal to DistTrain. They overlook the model and data heterogeneity in multimodal LLMs. DistTrain also integrates several of these optimizations in training the LLM backbone.
Multimodal model training. Small multimodal models (e.g., CLIP [61] and LiT [62]) have been widely studied in recent years. Many system optimizations have been proposed to train such multimodal models efficiently. DistMM [63] tackles model heterogeneity by introducing modality-aware placement and partitioning to evenly distribute workload. GraphPipe [64] presents graph pipeline parallelism, addressing graph dependencies in multimodal models to minimize pipeline bubbles. Yet, these advancements primarily enhance small multimodal models trained on tens of GPUs. They fall short for scaling up to meet the demands of integrating multimodal models with LLMs which necessitate training across thousands of GPUs. This gap underpins the motivation behind DistTrain, designed to meet the unique challenges of multimodal large language model training.
Multimodal LLM serving. LLM serving has been widely studied in recent years. Orca [65], FastServe [66], and VTC [67] propose iteration-level scheduling algorithms to improve the serving quality. DistServe [68] and Splitwise [69] propose disaggregated serving for prefill and decoding to improve the serving throughput. vLLM [70], RAGCache [71], and SGLang [72] propose prefix caching to reuse the KV tensors. However, the serving of multimodal LLMs remains under-explored. DistTrain’s core insights (e.g., disaggregation) are applicable to multimodal LLMs serving. Our future work will delve into the specific challenges posed by multimodal LLMs serving systems.
11 Conclusion
We present DistTrain, an efficient and adaptive framework to reform the training of multimodal LLMs. We identify the key challenges in training multimodal LLMs, i.e., model heterogeneity and data heterogeneity. DistTrain introduces disaggregated training to address these challenges, including disaggregated model orchestration to address model heterogeneity and disaggregated data reordering to address data heterogeneity. We evaluate DistTrain on production cluster with thousands of GPUs and show that it achieves 54.7% MFU and outperforms Megatron-LM by up to 2.2 on throughput.

### v3

@@ 7.1. Overall Performance
We first compare the overall performance of DistTrain against Megatron-LM on a large GPU cluster (up to 1296 GPUs). We retrofit Megatron-LM to support multimodal LLM training by integrating modality encoder and generator into the training pipeline. Megatron-LM employs monolithic model orchestration as described in §2.1. In Megatron-LM, we set the PP size of the LLM backbone to 1, 2, and 10 for Llama3-7B, Llama3-13B, and Llama3-70B, respectively. PP size is set to 1 for modality encoder and generator. We set TP size to 8, as each node consists of 8 GPUs connected by high-bandwidth NVLink. As for DistTrain, the parallelism strategy is determined by disaggregated model orchestration. In our experiments, one GPU is able to facilitate training ViT and SD. We replicate the modality encoder and generator across the GPUs within the TP group to process different images, whereas TP itself is not used. We set the global batch size to 1920.
The experimental results are shown in Figure 13 and Figure 14. Figure 13 shows the MFU. Figure 14 shows the training throughput. Due to the different model orchestration strategies of DistTrain and Megatron-LM, DistTrain uses 1056, 1216, and 1176 GPUs for MLLM-9B, MLLM-15B, and MLLM-72B, respectively. Megatron-LM uses 1296, 1280, and 1152 GPUs, respectively. Notably, within a total budget of 1296 GPUs, DistTrain intentionally allocates fewer resources in some cases because adding more GPUs yields no further improvements in training throughput. This resource efficiency frees the remaining GPUs for concurrent tasks such as fine-tuning or inference. We summarize the findings as follows.
@@
As shown in Figure 13, DistTrain achieves 51.8%-54.7% MFU in large-scale multimodal LLM training. This performance closely approximates that of state-of-the-art unimodal (i.e., text) LLM training (Jiang et al., 2024), which demonstrates the effectiveness of DistTrain in addressing the model and data heterogeneity in multimodal LLM training.
@@
DistTrain significantly outperforms Megatron-LM, delivering 1.7–2.8 the MFU and 1.7–2.2 the training throughput when training MLLM-9B and MLLM-15B with a similar number of GPUs. These performance gains largely stem from DistTrain’s disaggregated model orchestration. Megatron-LM’s monolithic strategy often leads to GPU underutilization, since it assigns too many GPUs to the modality encoder and generator. In contrast, DistTrain adaptively adjusts model orchestration based on specific model and data demands. Additionally, DistTrain’s disaggregated data preprocessing technique further improves efficiency.
@@
In the MLLM-72B training scenario, DistTrain outperforms Megatron-LM by 1.2 on MFU and 1.3 on training throughput with a similar number of GPUs. The high image resolution increases the execution time of the multimodal module, which introduces pipeline bubbles in LLM backbone. DistTrain addresses this by allocating additional GPUs to these modules to balance the pipeline. The disaggregated data preprocessing strategy continues to mitigate data heterogeneity, thereby increasing training efficiency.
7.2. Ablation Study
@@
In this subsection, we perform microbenchmarks to evaluate the effectiveness of each DistTrain’s technique and utilize up to 96 GPUs. We reduce the cluster scale due to our limited budget. We set the global batch size of training to 128, 64, and 40 for MLLM-9B, MLLM-15B, and MLLM-72B, respectively.
@@
Disaggregated model orchestration. We evaluate the MFU and training throughput using the model orchestration strategies of DistTrain, Megatron-LM, and DistMM∗. DistMM∗ allocates GPUs according to the computation demands (flops) of each module. DistTrain uses 96, 80, and 82 GPUs for MLLM-9B, MLLM-15B, and MLLM-72B, respectively. Megatron-LM uses 96, 96, and 96 GPUs, while DistMM∗ uses 88, 76, and 90 GPUs. The experimental results are shown in Figure 15. DistTrain consistently outperforms the baseline strategies, achieving 1.3–2.7 higher MFU and 1.4–2.7 higher training throughput. Although DistMM∗ outperforms Megatron-LM, it still lags behind DistTrain since it neglects the intricate performance model (§4.2) of multimodal LLM training. DistTrain’s disaggregated model orchestration optimally balances computational loads across the three modules and achieves high resource utilization. In production training, the running time of disaggregated model orchestration is less than one second, negligible compared to the days required for full training.
@@
We also evaluate the running time of DistTrain’s disaggregated model orchestration to determine the optimal resource and parallelism configurations under different training settings, as detailed in Table 3. The algorithm completes in under one second. The overhead is negligible compared to the days or even weeks required for overall training.
Disaggregated data preprocessing. We evaluate the effectiveness of DistTrain’s disaggregated data preprocessing by comparing it against Megatron-LM’s data preprocessing, while keeping other components the same. Megatron-LM’s data preprocessing uses random ordering for training data. The effectiveness is gauged through metrics such as MFU and training throughput. We use the optimal resource allocation and parallelism strategy decided by DistTrain’s disaggregated model orchestration. Given that the model orchestration strategy remains unchanged, we do not indicate the number of GPUs. The experimental settings are the same as those in the experiment to evaluate disaggregated model orchestration. The results are shown in Figure 16. DistTrain consistently outperforms the baseline, achieving 1.03-1.11 higher MFU and throughput. The performance gap becomes more pronounced as the model size decreases. This is because the smaller model size leads to a higher data parallelism (DP) size, which causes more intra-microbatch heterogeneity. In essence, DistTrain’s two-level data reordering effectively mitigates data heterogeneity and enhances training efficiency. We do not measure the running time of the reordering algorithm as it operates on dedicated CPU nodes asynchronously. The data preprocessing overhead with and without disaggregation is evaluated in §7.3.
@@
7.3. Case Study
In this subsection, we first evaluate the overhead of data preprocessing. We then evaluate DistTrain under different frozen training settings.
@@ 7.3. Case Study
Overhead of data preprocessing. We conduct an experiment to evaluate the overhead of data preprocessing, including decompression and reordering. Setting the DP size to one, we measure the average data preprocessing time per iteration on the GPU training side. We then compare data preprocessing time with and without disaggregated data preprocessing and use varying numbers of images and different image resolutions for one training iteration. The results, depicted in Figure 17, indicate disaggregating data preprocessing from training reduces preprocessing time from seconds to milliseconds. The first parameter in the x-axis represents the number of images, while the second parameter denotes the image resolution. In production training (§7.1), iteration times range from seconds to tens of seconds. Without disaggregation, preprocessing overhead is in seconds, significantly interferes with training. With disaggregated data preprocessing, the overhead reduces to milliseconds, which is negligible relative to total iteration time.
(a) All modules freezing. (b) Encoder-only training. (c) LLM-only training. (d) Generator-only training.
(a) All modules freezing. (b) Encoder-only training. (c) LLM-only training. (d) Generator-only training.
Frozen training. In real production training, some modules in multimodal LLM are frozen to stabilize training loss and enhance model effectiveness during different training phases (Yin et al., 2023). We conduct a frozen training experiment under four specific training settings: complete module freezing (i.e., training projectors only), encoder-only training, LLM-only training, and generator-only training. In these scenarios, frozen modules neither compute weight gradients in backward pass nor update weights. The frozen modules still perform computations in forward pass. All other experimental setup aligns with those detailed in §7.2. DistTrain uses 96, 80, and 82 GPUs for MLLM-9B, MLLM-15B, and MLLM-72B under all three frozen settings. Megatron-LM uses 96, 96, and 96 GPUs for MLLM-9B, MLLM-15B, and MLLM-72B under all three frozen settings. The experimental results are in Figure 18 and Figure 19. DistTrain consistently outperforms Megatron-LM across all frozen training configurations, achieving 1.4–2.9 higher MFU and 1.2–2.9 higher training throughput. This pronounced performance gap underscores the challenges posed by Megatron-LM’s monolithic model orchestration in more complex training environments. In contrast, DistTrain adaptively adjusts model orchestration based on training settings and achieves high resource utilization.
8. Discussion
@@
Parallelism strategies. Many studies (Zheng et al., 2022; Wang et al., 2019; Jia et al., 2019) optimize parallelism strategies to accelerate deep learning model training. However, they fall short of holistically considering data, pipeline, and tensor parallelism in multimodal LLM training, due to the large search spaces in large clusters. This limitation results in inefficient and sub-optimal parallelism strategy. Additionally, these methods generally assume homogeneity of training data—all training samples consume the same amount of computation. This assumption does not hold for multimodal LLMs where training samples of different modalities vary significantly in volume and computation demand. DistTrain leverages the specific training pattern of multimodal LLMs, and formulates a model orchestration problem for multimodal LLMs that integrates tensor, pipeline, and data parallelism simultaneously. Besides, DistTrain leverages disaggregated data reordering to address data heterogeneity.
@@
Sequence and expert parallelism. Sequence parallelism (SP) (Li et al., 2021) is designed to partition the training sequence into multiple subsequences for parallel training. It addresses the challenges of processing long sequences in LLMs. Expert parallelism (EP) (Liu et al., 2023b), specifically devised for mixture-of-experts (MoE) LLMs (Du et al., 2022), enables parallel training of multiple feedforward network (FFN) experts. These parallelism strategies are orthogonal to multimodal LLM training. In DistTrain, both SP and EP are integrated into the LLM backbone training.
Heterogeneous hardware. By disaggregating three modules (i.e., the modality encoder, LLM backbone, and generator) for multimodal LLM training, DistTrain supports using heterogeneous hardware for different modules to achieve varius goals like improving performance, reducing cost and increasing energy efficiency. This strategy assigns distinct computational workloads to the most appropriate hardware. For instance, ViT encoder is significantly less compute-intensive than LLM backbone. Consequently, we can place ViT encoder on more economical GPUs (e.g., NVIDIA L20). Such benefits are demonstrated in other LLM systems such as MegaScale-Infer (Zhu et al., 2025) and Splitwise (Patel et al., 2024).
Disaggregated data preprocessing. The principal technique in DistTrain’s disaggregated data preprocessing is data reordering that addresses training stragglers caused by data heterogeneity. Although data reordering does not inherently necessitate disaggregation, we adopt disaggregation to mitigate the preprocessing overhead and improve system elasticity (e.g., allocating any arbitrary number of CPUs for preprocessing tasks). This approach is analogous to resource disaggregation in other systems scenarios, which aims to improve system elasticity and efficiency (Shan et al., 2018; Wang et al., 2022a).
9. Related Work
LLM training. Many efforts have been made to optimize LLM training. For LLM pre-training, Megatron-LM (Shoeybi et al., 2019) and DeepSpeed-Megatron (Smith et al., 2022) propose custom 3D-parallelism strategies. DeepSpeed-ZeRO (Rajbhandari et al., 2020) and Pytorch-FSDP (Zhao et al., 2023) reduce redundant memory consumption in data parallelism. A set of works (Chen et al., 2024; Jiang et al., 2024; Chang et al., 2024; TE, 2024; Wang et al., 2022b) overlap the communication and computation operators in LLM training. These systems optimizations of LLM training are orthogonal to DistTrain. They do not address model and data heterogeneity in multimodal LLM training. DistTrain integrates several of these optimizations in training the LLM backbone.
Traditional multimodal training. Traditional multimodal models (e.g., CLIP (Radford et al., 2021) and LiT (Zhai et al., 2022)) have been widely studied in recent years. Many systems optimizations have been proposed to train such multimodal models efficiently. DistMM (Huang et al., 2024) optimizes multiple parallel encoders training for contrastive learning. Yet, these advancements primarily enhance traditional multimodal models (e.g., CLIP (Radford et al., 2021) and LiT (Zhai et al., 2022)), which do not integrate LLM for understanding and generation (e.g., Flamingo (Alayrac et al., 2022) and LlaVa (Liu et al., 2024a)). They do not support encoder, LLM backbone and generator training pipeline in multimodal LLMs. Therefore, the state-of-the-art solution is to leverage Megatron-LM to train multimodal LLMs at large scale.
10. Conclusion
We present DistTrain, a disaggregated multimodal LLM training system. We identify the key challenges in training multimodal LLMs, i.e., model heterogeneity and data heterogeneity. DistTrain introduces disaggregated model orchestration to address model heterogeneity and disaggregated data preprocessing to address data heterogeneity. We evaluate DistTrain on a large production cluster with thousands of GPUs and show that it achieves 54.7% MFU and outperforms Megatron-LM by up to 2.2 on training throughput.
This work does not raise any ethical issues.
Acknowledgments. We sincerely thank the anonymous reviewers for their valuable feedback on this paper. Xin Jin is the corresponding author. Zili Zhang, Yinmin Zhong, and Xin Jin are also with the Key Laboratory of High Confidence Software Technologies (Peking University), Ministry of Education.
This work was supported in part by the National Key Research and Development Program of China under Grant 2022YFB4500700, the Scientific Research Innovation Capability Support Project for Young Faculty under Grant ZYGXQNJSKYCXNLZCXM-I1, the Fundamental Research Funds for the Central Universities, Peking University, and the National Natural Science Foundation of China under Grant 62172008.
Appendices are supporting material that has not been peer-reviewed.
Appendix A Appendix
A.1. Mitigating TP overhead with StepCCL
Tensor Parallelism (TP) is commonly adopted to facilitate training large Transformer-based models with multiple GPUs connected with high bandwidth (e.g., NVLinks). Specifically, TP divides the linear layers, into smaller sub-modules, which are then distributed across the GPUs. After parallel computation, all GPUs perform collective communication to aggregate data and produce identical results. The TP communication overhead severely degrades overall performance.
We implement the communication overlap with an in-house collective communication library called StepCCL to reduce the TP overhead. StepCCL is a PyTorch custom plugin that performs cross-GPU collective communication, including allgather, reduce-scatter, and allreduce, which is similar to NCCL. However, NCCL occupies several CUDA Streaming Multiprocessors (SMs) for executing its communication kernel and is known to harm the performance of its concurrent GEMM (matrix multiplication) (Rashidi et al., 2021). To solve this, StepCCL leverages the DMA engine directly to perform data transmission without using any SM at all. This enables StepCCL and GEMM to run simultaneously on a GPU without slowing down each other. This cornerstone facilitates our subsequent development of communication overlap.
Figure 20 shows an example of how StepCCL works in overlapping the allgather (AG) operation with GEMM. We start by decomposing the single GEMM and the corresponding communication into several smaller pairs. Each small communication operation starts sequentially on a communication stream, with its paired GEMM executed on the default computation stream. The communication overhead is fully hidden except for the first allgather.111If the number of allgather/GEMM is large enough, the only allgather in the critical path should have negligible overhead. But dividing a large GEMM into finer granularity sometimes could lead to overall slowdown. In practice, the number is actually configurable. After all GEMMs finish, we perform an extra layout remapping operation (usually with negligible overhead) to ensure identical results with the baseline. Figure 21 describes the details of the layout remap process. In some rare cases during backward propagation, we find the remap overhead is high due to certain model dimensions. To mitigate this, we further overlap the remap with the computation of the weight gradients, so eventually we nearly get the full performance gain of the communication overlap.
Finally, although the overlap idea is also studied in many related works (Chang et al., 2024; TE, 2024; Wang et al., 2022b), we highlight the key differences of ours. Unlike prior work that fuses GEMM with TP communication into one CUDA kernel (Chang et al., 2024; TE, 2024), we choose a modular design and do not use fusion for more flexibility. For example, when TP communication is longer than GEMM, fusing them cannot fully hide the communication overhead. However, with the modular design, we are able to hide the communication with other modules without dependency (e.g., in cross-attention), which is not possible with the fused implementation. This enables broader adoption of StepCCL in many other scenarios.
Evaluation. To evaluate the effectiveness of StepCCL in mitigating the TP overhead, we conduct an experiment that measures the iteration time of the LLM backbone with training of one single PP stage (i.e., one minimal TP group) under various TP sizes. We compare the iteration time with and without StepCCL enabled. The results are shown in Figure 22. StepCCL significantly reduces the iteration time by overlapping the TP communication with computation. It outperforms the baseline by 1.1-1.12 when the TP size is 4 and 1.15-1.17 when the TP size is 8. The gains are more pronounced at large TP size, where communication overhead is more substantial. These findings confirm that StepCCL effectively mitigates TP overhead.